OpenStudy (anonymous):

dy/dx of xy^2+4y-5x^2 (implicit differentation) plz help and show work !!!

zepdrix (zepdrix):

There is no equality sign.... xy^2+4y-5x^2 ... = 0 ?

OpenStudy (anonymous):

ya i think so

OpenStudy (anonymous):

Basically for implicit differentiation, every time you take the derivative of y add a dy/dx after it. Ex) 3x^2 + 2y^3 = 0 6x + 6y^2 dy/dx = 0 Now we have to solve for dy/dx though 6y^2 dy/dx = -6x dy/dx = -6x/6y^2 dy/dx = -x/y^2 Did this help?

OpenStudy (anonymous):

sorta..not really tho...

OpenStudy (anonymous):

What specifically are you confused about?